# CrEIMBO: Cross Ensemble Interactions in Multi-view Brain Observations

### v1

@@ 4.2 CrEIMBO Discovers Multi-Regional Dynamics in Human Electrophysiology Data
CrEIMBO offers several advantages over other models of whole-brain multi-session data. Chiefly, it identifies a cross-session shared latent space where non-stationary ensemble interactions are governed by a set of global sub-circuits. By structuring the time-varying dynamics in terms of a basis set of sub-circuits, CrEIMBO allows for the exploration of active sub-circuits across sessions and aiding in the identification of subject-to-subject and cross-session variability in neural activity. Additionally, constraints on the loading matrices $\bm{A}$ enable the identification of area-specific ensembles, facilitating the understanding of within-area compared to between-area interactions. Moreover, the simultaneous fitting of the ensembles and dynamics automatically aligns cross-session ensembles in terms of functionality, which is critical for studying cross-session variability. We demonstrated the efficacy of CrEIMBO through both synthetic experiments and neural recordings. On synthetic data, CrEIMBO consistently recovered ground truth components with high accuracy. In our neural recordings analysis, CrEIMBO found cross-regional motifs that span interactions across sessions and subjects.
Limitations and future work: While CrEIMBO leverages information from multiple brain sessions, the most modern Electrophysiology tools are still far from observing every neuron or region simultaneously. Consequently, and as with all dynamical systems models, obscured interactions with missing (i.e., unrecorded) neurons within the brain may limit the precision of the identified interactions. Additionally, CrEIMBO’s current implementation does not include an offset term. Adding this would increase complexity and require thorough exploration; hence, we will address it in future work. We limit CrEIMBO here to linear projections from the ensemble-space to per-session observation space. Extending this aspect of the model to more flexible, non-linear projections while maintaining interpretability is challenging and a future goal.
CrEIMBO’s effectiveness in aligning ensembles across sessions relies on the assumption that these sessions share commonalities in their active sub-circuits. If sessions operate with entirely different sub-circuits, cross-session ensemble alignment is not guaranteed. Another limitation is that we assume Gaussian observation statistics, i.e., spiking data must be converted to rates to use CrEIMBO. Future research could incorporate non-Gaussian likelihoods, such as Poisson for spike counts. Finally, CrEIMBO rests on the mathematical foundation of dictionary learning, which can be computationally intensive at higher dimensions. To enhance computational efficiency, we plan to implement parallel processing, data batching, and GPU processing in future iterations.
We also look to leverage CrEIMBO to identify pathological or cognitive states during which specific sub-circuits manifest exclusively (e.g., neural circuitry changes during epileptic attacks). Another extension for the current method includes extending the global cross-session circuits set to enable small, controlled structural adaptations of the sub-circuits between sessions, e.g., through an $\ell_{2}$ regularization on the distance between cross-session corresponding sub-circuits $\|\bm{f}_{k}^{d}-\bm{f}_{k}^{d^{\prime}}\|_{F}^{2}$.
Acknowledgments
N.M. was supported by the Kavli Discovery Award of the Kavli Foundation.
@@ Appendix A Notations
Appendix B Differences from dLDS
While our proposed model draws inspiration from dLDS [16], it significantly diverges in its approach, particularly in ensemble identification and the details thereof. Below we list the key differences:
  1. 1.
    Incorporation of sparsity constraints on the operators $f$.
  2. 2.
    Extension to multi-sessions through shared $f$s and distinct projections ($D$) instead of coefficients ($c$).
  3. 3.
    Imposition of regional structure on $D$.
  4. 4.
    Integration of graph-driven sparsity on $D$.
  5. 5.
    Modifications to the inference procedure inspired by linocs.
  6. 6.
    Ability to capture inter-subject variability.
  7. 7.
    Adaptation to handle variations in trial duration, fluctuations in neuron count, and changes in distributions between sessions.
Appendix C Differences from SiBBlInGS
Although our proposed model draws inspiration from SiBBlInGS [17] for ensemble identification, it introduces several fundamental differences:
  1. 1.
    SiBBlInGS does not consider interactions between ensembles, whereas our model incorporates ensemble interactions.
  2. 2.
    SiBBlInGS does not address differences in the number of neurons between subjects, whereas our model accounts for this variability.
  3. 3.
    SiBBlInGS lacks consideration for region-specific structure, focusing primarily on a single block of $D$, whereas our model imposes region-specific structure on $D$.
  4. 4.
    SiBBlInGS does not account for the evolution of the ensemble activity space, whereas our model incorporates this aspect.
Appendix D Parameters Range for Synthetic Data Experiment
The full set of parameters for the synthetic experiments are available in Table 1.
@@ Appendix D Parameters Range for Synthetic Data Experiment
Appendix E Parameters for human multi-regional experiment
@@
Appendix F Information about the baselines:
@@
    All Regions (sparse)”: Similar to the previous one but without applying sparsity on the ensemble matrix.

### v2

@@ 5 Discussion, Limitations, and Future Work
Here, we introduced CREIMBO—a novel approach for uncovering multi-regional dynamics in neural data collected across multiple sessions. CREIMBO addresses the challenge of integrating non-simultaneous neural recordings by joint dynamical inference and sparsity regularization to capture the underlying neural sub-circuits governing brain activity. We further demonstrated the efficacy of CREIMBO through multiple synthetic and neural data, and found that CREIMBO recovered ground truth components and is robust to noise in identifying cross-regional motifs that span cross-session interactions. CREIMBO offers several advantages over existing methods. Chiefly, it identifies a cross-session shared latent space where non-stationary ensemble interactions are governed by a time-varying decomposition of universal basis dynamics. By structuring these dynamics in terms of global sub-circuits, CREIMBO allows the discovery of sub-circuits meaningful for various cognitive processes, enabling the identification of variability in neural activity across subjects and sessions.
Limitations & Future Directions: An important feature of CREIMBO is its ability to unify sessions with different neuron subsets through the universal dictionary of dynamical interactions prior that aligns ensembles in terms of functionality. This can enable the inference of their activity even if they are not observed in some sessions. This ability, however, depends on the overlap between the used basis interactions across sessions, the distinctiveness of different $\bm{f}_{k}$s, and the premise that the basis interactions capturing interactions with ensembles missing from a session include dynamic values for ensembles observed in that same session (Sec. Assumptions). Another limitation is CREIMBO’s reliance on linear projections from ensemble to session observation space, restricting flexibility. Extending to non-linear projections offers potential for development, though it introduces computational and interpretability challenges. Finally, CREIMBO uses dictionary learning, which is computationally demanding, and future iterations will include parallel processing. Extending CREIMBO to additional data types and applications, including non-neural data (e.g., immune-cell counts), or for identifying pathologies, is an exciting future step. Moreover, integrating CREIMBO with mDLAG (Gokcen et al., 2024), with the latter identifying optimal communication delays and the former leveraging these to further identify the underlying set of co-active LDSs, presents an exciting future work.
@@ Appendix A Notations
Appendix B Synthetic Data Experiment Details
We simulated the synthetic data to represent the non-stationary dynamics of brain activity including multi-region interactions. Current assumptions about brain activity suggest Zeki (2015); Nelson & Bower (1990) it functions through distributed, parallel processing across multiple regions, potentially involving co-active circuits that can process a range of variables (e.g., task demands, feedback signals, and sensory information), simultaneously. This biological reality thus motivated our data generation approach, to reflects the multi-process nature of neural activity, where different processes can interact and contribute to brain dynamics. We simulated $D=5$ sessions, each sharing a common set of basis rotational dynamics, denoted as $\{\bm{f}_{k}\}_{k=1}^{K=3}$, with $p_{j}=3$ regions of interest. Different regions were associated with different ensembles, and the dynamics for each session were modeled using sparse decomposition of the basis dynamics. For each session, the dynamics’ coefficients, $\bm{c}_{t}^{k}$, were generated over $T=500$ time points. These coefficients were generated by initially giving them binary values (active vs. inactive states), with random switches occurring every $x$ time points, where $x$ was uniformly drawn between 10 and 30. The sparsity constraint was maintained at each time point, ensuring that only one dynamic was active at any time ($\|\bm{c}_{t}^{k}\|_{0}=1\forall t$).
Then, to introduce smooth transitions between switching states, a Gaussian convolution with a time window of 6 points (standard deviation = 0.7) was applied to the coefficients, to promote a smoother / more gradual changes between active states. Different sessions maintained the same sequence of active dynamics but allowed for varying durations of each state, introducing temporal variability across sessions.
Each session $d$ was augmented with a unique projection matrix $\bm{A}^{d}$ that captures the mapping between the latent space and the observations (neural) space in the respective regions. These projection matrices $\bm{A}^{k}$ consisted of $P=3$ columns, each corresponding to one of the three regions. The number of neurons in each region was randomly chosen between 4 and 9, with the count varying independently across regions and sessions. This variability ensured that different sessions were not identical in their neural configurations. Latent dynamics for each session were generated using the recursive equation $\bm{x}_{t}=\sum_{k=1}^{K}\bm{c}_{kt}\bm{f}_{k}\bm{x}_{t-1}$, starting from an initial state $\bm{x}_{0}=[1,-1,1]^{\top}$. CREIMBO was trained based on the observations only, which were defined for each session as $\bm{y}^{d}=\bm{A}^{d}\bm{x}^{d}+\epsilon$ where $\epsilon$ is an addition of i.i.d Gaussian noise with standard-deviation of $\sigma_{\text{noise}}=0.1$
The full set of parameters for the synthetic experiment are available in Table 1.
@@ Appendix B Synthetic Data Experiment Details
Appendix C Parameters for human multi-regional experiment
@@
Appendix D Information about the baselines:
We distinguish between two types of baselines: (1) ablation experiments and (2) comparison to the performance of entirely different methods. It is important to note that the latter cannot fully capture the unique capabilities of CREIMBO, particularly in identifying basis ensemble interactions shared across sessions through sparse decomposition. Therefore, we only compare certain aspects that can be extracted from other methods (e.g., multi-regional SLDS, which requires the same number of neurons per session). However, these comparisons do not imply that other models outperform or underperform CREIMBO; they simply offer different insights, with varying strengths and limitations depending on the application.
Importantly, since the sub-circuits $\{\bm{f}_{k}\}$ is invariant to its ordering, we used SciPy’s implementation of the ‘linear_sum_assignment‘ problem Crouse (2016) to re-order the identified sub-circuits before comparing them with the ground truth, matching the $\bm{f}$s values by minimizing the $\ell_{2}$ error. Ablation Experiments:
@@ Appendix D Information about the baselines:
    All Regions (non-sparse)”: Similar to the previous one but without applying sparsity on the ensemble matrix.